Gauge length prior to starting =50mm



Length after fracture =66.4mm



0.1% proof load=18.5kn



Maximum load=26.6kn



Original diameter =10mm





Calculate using the above data the



2. 0.1 proof stress



3. Tensile strength

1
Expert's answer
2021-11-29T11:48:01-0500

1.

0.1% proof stress means 0.1% of E (Young's modulus of elasticity):

"E=\sigma/\varepsilon"


stress:

"\sigma=F/A"

where F is 0.1% proof load,

A is cross-sectional area

"\sigma=\frac{18.5\cdot10^3}{\pi(10\cdot10^{-3})^2}=5.89\cdot10^7" Pa


stain:

"\varepsilon=x/L"

where x is extension,

L is length


"\varepsilon=\frac{66.4-50}{50}=0.328"


"E=5.89\cdot10^7/0.328=1.796\cdot10^8" Pa


2.

Tensile strength:

"F_{max}/A=\frac{26.6\cdot10^3}{\pi(10\cdot10^{-3})^2}=8.47\cdot10^7" Pa